ram is 5 years elder than Lakshman who is 4 years elder than Bharath. The sum of their ages is 61. Find the age of Lakshman?
Answer by Alan3354(69443)

ram is 5 years elder than Lakshman who is 4 years elder than Bharath. The sum of their ages is 61. Find the age of Lakshman?
============
R = L + 5
L = B + 4
--> R = B + 9
----
R + B + L = 61
--
B+9 + B+4 + B = 61
3B = 48
B = 16
L = 20